# On Rollouts in Model-Based Reinforcement Learning

### v1

@@ 4.1 Theoretical Setup
$\mathbb{E}\left[\hat{S}_{t+1}|S_{t},A_{t}\right]=\mathbb{E}\left[\check{S}_{t+1}|S_{t},A_{t}\right]\quad\forall(S_{t},A_{t})\in\mathcal{E}.$ (10)
@@ 4.2 Decomposing the model state in signal and noise
concerning the $\check{S}_{t+1}$ and $\Delta_{t}$ represented by $\mu^{\Delta}(S_{t},A_{t})$ the model bias, $\Sigma^{\Delta}(S_{t},A_{t})$ the epistemic variance with Cholesky decomposition $L^{\Delta}(S_{t},A_{t})$, and $N_{t}$ the epistemic noise.
@@ 4.4 Rollout Termination Criteria
First, Infoprop rollouts build on the assumption that mudel usage is restricted to a sufficiently accurate subset $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{A}$, following the ideas of Frauenknecht et al. (2024).
@@ 6.3 Performance Evaluation
A major concern of this work is the consistency of model-based rollouts with the environment distribution. Figure 5(b) depicts the data distribution in $\mathcal{D}_{\mathrm{env}}$ and $\mathcal{D}_{\mathrm{mod}}$ of the respective Dyna-style approaches throughout training for the $11^{\text{th}}$ dimension of the hopper state. The figure shows a histogram over state values over the course of training. It can be seen that the model data distribution of Infoprop-Dyna closely follows the distribution observed in the environment, while both the data from MBPO and MACURA show severe outliers. This is the case, even though the rollout data in Infoprop-Dyna is obtained from substantially longer rollouts as can be seen from Figure 5(a) which indicates the capabilities of the Infoprop rollout mechanism.
@@ 6.4 Limitations and Outlook
Despite the excellent quality of model-generated data with the Infoprop rollout, the limitations of Infoprop-Dyna are most apparent on MuJoCo humanoid with results provided in E.3. These show instabilities in learning and point to structural problems when integrating Infoprop rollouts naively into standard Dyna-style architectures Janner et al. (2019).
Figure 5(b) shows that the long rollouts of Infoprop-Dyna can cause rapid distribution shifts in $\mathcal{D}_{\mathrm{mod}}$, especially early in training. These nonstationary buffers are a challenge to deep Q-learning methods Mnih et al. (2015). Another issue is primacy bias in model learning Qiao et al. (2023), where the model overfits to initial data and subsequently struggles to generalize, as seen in the decreasing rollout length for the ant environment in Figure 5(a). The main problem with Infoprop-Dyna is likely overfitting critics and plasticity loss Nikishin et al. (2022); D’Oro et al. (2023), as also reported by Frauenknecht et al. (2024) for Dyna-style MBRL trained on high-quality data. We provide an ablation on this observation and sketch methods to counteract this phenomenon in Appendix E.4.
@@ Acknowledgments
We thank Christian Fiedler, Pierre-François Massiani, and David Stenger for the fruitful discussions on the work presented in this paper. This work is funded in part under the Excellence Strategy of the Federal Government and the Länder (G:(DE-82)EXS-SF-SFDdM035) and the German Federal Ministry of Education and Research (BMBF) under the Robotics Institute Germany (RIG), which the authors gratefully acknowledge. Further, the authors gratefully acknowledge the computing time provided to them at the NHR Center NHR4CES at RWTH Aachen University (project number p0022301). This is funded by the Federal Ministry of Education and Research, and the state governments participating on the basis of the resolutions of the GWK for national high performance computing at universities (www.nhr-verein.de/unsere-partner).

### v2

@@ 4.1 Theoretical Setup
$\mathbb{E}[\hat{S}_{t+1}|S_{t},A_{t}]=\mathbb{E}\left[\check{S}_{t+1}|S_{t},A_{t}\right]\quad\forall(S_{t},A_{t})\in\mathcal{E}.$ (10)
@@ 4.2 Decomposing the model state in signal and noise
concerning $\check{S}_{t+1}$ and the model error $\Delta_{t}$ represented by $\mu^{\Delta}(S_{t},A_{t})$ the model bias, $\Sigma^{\Delta}(S_{t},A_{t})$ the epistemic variance with Cholesky decomposition $L^{\Delta}(S_{t},A_{t})$, and $N_{t}$ the epistemic noise.
@@ 4.4 Rollout Termination Criteria
First, Infoprop rollouts build on the assumption that model usage is restricted to a sufficiently accurate subset $\mathcal{E}\subseteq\mathcal{S}\times\mathcal{A}$, following the ideas of Frauenknecht et al. (2024).
@@ 6.3 Performance Evaluation
A major concern of this work is the consistency of model-based rollouts with the environment distribution. Figure 5(b) depicts the data distribution in $\mathcal{D}_{\mathrm{env}}$ and $\mathcal{D}_{\mathrm{mod}}$ of the respective Dyna-style approaches throughout training for the $11^{\text{th}}$ dimension of the hopper state. The distributions are illustrated via histograms over environment steps. It can be seen that the model data distribution of Infoprop-Dyna closely follows the distribution observed in the environment, while both the data from MBPO and MACURA show severe outliers. This is the case, even though the rollout data in Infoprop-Dyna is obtained from substantially longer rollouts as can be seen from Figure 5(a) which indicates the capabilities of the Infoprop rollout mechanism.
@@ 6.4 Limitations and Outlook
Despite the excellent quality of model-generated data with the Infoprop rollout, the limitations of Infoprop-Dyna are most apparent on MuJoCo humanoid with results provided in Appendix E.3. These show instabilities in learning and point to structural problems when integrating Infoprop rollouts naively into standard Dyna-style architectures Janner et al. (2019).
Figure 5(b) shows that the long rollouts of Infoprop-Dyna can cause rapid distribution shifts in $\mathcal{D}_{\mathrm{mod}}$, especially early in training. These nonstationary buffers are a well-known challenge to deep Q-learning methods Mnih et al. (2015). Another issue is primacy bias in model learning Qiao et al. (2023), where the model overfits to initial data and subsequently struggles to generalize, as seen in the decreasing rollout length for ant in Figure 5(a). The main problem with Infoprop-Dyna is likely overfitting critics and plasticity loss Nikishin et al. (2022); D’Oro et al. (2023), as also reported by Frauenknecht et al. (2024) for Dyna-style MBRL trained on high-quality data. We provide an ablation on this observation and sketch methods to counteract this phenomenon in Appendix E.4.
@@ Acknowledgments
We thank Christian Fiedler, Pierre-François Massiani, and David Stenger for the fruitful discussions on the work presented in this paper. This work is funded in part under the Excellence Strategy of the Federal Government and the Länder (G:(DE-82)EXS-SF-OPSF854) and the German Federal Ministry of Education and Research (BMBF) under the Robotics Institute Germany (RIG), which the authors gratefully acknowledge. Friedrich Solowjow is supported by the KI-Starter grant by the state of NRW. Further, the authors gratefully acknowledge the computing time provided to them at the NHR Center NHR4CES at RWTH Aachen University (project number p0022301). This is funded by the Federal Ministry of Education and Research, and the state governments participating on the basis of the resolutions of the GWK for national high performance computing at universities (www.nhr-verein.de/unsere-partner).